OpenStudy (anonymous):

Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)? 325 9750 355 295

OpenStudy (anonymous):

@Dimples

OpenStudy (anonymous):

Okay you have two functions here right? Function f(x) which is 13x2 or 13 x (x) x 2 (or should that be ^2?). and Function g(x) which is 7x - 5. The problem wants you to solve (f+g)(x) or (f+g)(5) So just substitute x for 5 and solve both function then add them together.

OpenStudy (anonymous):

so it would be etheir a b or d?

OpenStudy (anonymous):

BTW \[f(x) 13x^{2}\] And \[g(x)7x-5\]

OpenStudy (anonymous):

so it would be C

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

i need help on anther one

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of (f − g)(−2)? −8 −5 −11 3

OpenStudy (anonymous):

Okay this is the same thing as last time only x=-2 this time.

OpenStudy (anonymous):

so its -11

OpenStudy (anonymous):

dimples?

OpenStudy (anonymous):

I'm coming up with -13 I'm sure it's because I'm either misreading the original problem or you mistyped it, but if you solved and came up with -11 that I would venture that you are correct.

OpenStudy (anonymous):

so the awnser is -11?

OpenStudy (anonymous):

I cannot verify that because I cannot properly read the original problem. Can you enter the problem using the equation generator button in the lower left there?

OpenStudy (anonymous):

I've tried various versions of the problem, but cannot come up with any of the given answers.

OpenStudy (anonymous):

okay can you check my work and see if i did them right?

OpenStudy (anonymous):

As long as I can read the problem I can check them for you no problem.

OpenStudy (anonymous):

1.Choose the correct procedure for calculating (f − g)(−2) if f(x) = 13x2 and g(x) = 7x − 5. Calculate f(−2) and g(−2), then subtract them. Calculate f(−2) and g(−2), then add them together. Calculate f(−2) and g(−2), then write the results as an ordered pair. Calculate g(−2), plug the result into f(x) and simplify. i got B

OpenStudy (anonymous):

2. Choose the best description of how to calculate f[g(3)] if f(x) = 3x − 2 and g(x) = x2 − 1. Calculate f(3) and g(3), then multiply them together. Calculate f(3) and g(3), then write the results as an ordered pair. Calculate f(3) and g(3), then subtract them. Calculate g(3), plug the result into f(x) and simplify. i got A

OpenStudy (anonymous):

Wait one at a time.

OpenStudy (anonymous):

#1 is incorrect.

OpenStudy (anonymous):

okay are 1 and 2 right?

OpenStudy (anonymous):

how is 1 wrong?

OpenStudy (anonymous):

Both are wrong.

OpenStudy (anonymous):

Number 1 is wrong because you the answer you selected is performing the wrong operation.

OpenStudy (anonymous):

Look carefully at the problem. It is (f - g).

OpenStudy (anonymous):

so its subtract so its a?

OpenStudy (anonymous):

Number 2 is wrong because the answer you selected is performing the operation in the wrong order. And yes you are correct now #1 is A. :)

OpenStudy (anonymous):

is number 2 c?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

nvm number 2 is a?

OpenStudy (anonymous):

cause its multiplying

OpenStudy (anonymous):

In order to figure out f[g(3)], you have to first know what g(3) is.

OpenStudy (anonymous):

No # 2 is not a either.

OpenStudy (anonymous):

so its d?

OpenStudy (anonymous):

In that problem you are essentially figuring out a function of a function. Do you follow what I'm saying?

OpenStudy (anonymous):

yes so it is B!

OpenStudy (anonymous):

To calculate f(x) when x=g(3) then you have to first calculate g(3).

OpenStudy (anonymous):

No it is D. lol You have to know what g(3) is before you can solve f[g(3)]

OpenStudy (anonymous):

3. Consider the functions f(x) = 3x2 and g(x) = −2x − 5. What is the value of f[g(−4)]? 9 −31 27 3

OpenStudy (anonymous):

i put 27

OpenStudy (anonymous):

This is just like the last one. You have to calculate one first before you can calculate the other.

OpenStudy (anonymous):

Give me one moment and I'll calculate to check your answer.

OpenStudy (anonymous):

Yes 27

OpenStudy (anonymous):

4 . Consider the functions f(x) = 13x2 and g(x) = 7x − 5. What is the value of (f + g)(5)? 325 9750 355 295

OpenStudy (anonymous):

my awnser is 355

OpenStudy (anonymous):

Is f(x)\[13x^{2}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Then yes 355

OpenStudy (anonymous):

5. Choose the correct procedure for calculating (f + g)(−4) if f(x) = 2x + 1 and g(x) = x2 − 10. Calculate f(−4) and g(−4), then write the results as an ordered pair. Calculate f(−4) and g(−4), then add them together. Calculate f(−4) and g(−4), then subtract them. Calculate g(−4), plug the result into f(x) and simplify. i put b

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Is that all of them?

OpenStudy (anonymous):

no 3 more

OpenStudy (anonymous):

K

OpenStudy (anonymous):

Consider the functions f(x) = 8x + 2 and g(x) = 5x − 3. What is the value of ( f over g )(4)? Please show all of your work for full credit.

OpenStudy (anonymous):

my awnser : 8x+2/5x-3 then i pluged in 4 for x: 8(4)+2/5(4)-3= 32+2/20-3= 34/17= 2 final answer is 2

OpenStudy (anonymous):

Okay give me a minute on this one.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x)=8x+2\] \[g(x)=5x-3\] \[\left( \frac{ f }{ g } \right)\left( 4 \right)\] \[f(4)=8\times4+2=34\] \[g(4)=5\times4-3=17\] \[\frac{ 34 }{ 17 }=2\] Sorry it took so long I wanted to be sure that you could read all the work properly and yes you were correct.

OpenStudy (anonymous):

I think you just have 2 more?

OpenStudy (anonymous):

If you want me to check the last two, just let me know. Now that you saw the work on that last one I won't need to show you again and I'll be faster checking the answer.